# Enhancing Sensitivity for Di-Higgs Boson Searches Using Anomaly Detection and Supervised Machine Learning Techniques

### v1

@@ 4 Conclusion
In summary, we tested two fundamentally different ML approaches: one, a supervised classifier that requires both SM background and BSM signal events as input, and the other, an autoencoder-based anomaly detection approach that only requires SM events for training. The benchmarks were based on models with $X\rightarrow HH$ and $X\rightarrow SH$ boson decays, with the final state involving two leptons and two $b$-jets. The same input variables were used in both ML approaches, and the architecture of the hidden layers was identical. Therefore, these studies provide a compelling comparison of two distinct approaches that use the same inputs.
We concluded that both ML techniques were able to improve the significance of signal observations. The event classifier outperforms the autoencoder for the BSM events used in training, especially for low masses of the scalar $X$. This is unsurprising for the $HH$ signals, which were used to train the SC. However, it was unexpected to observe that the classifier works quite well even for the $SH$ signal events, which were not used in training. Its performance is comparable to that of the autoencoder. The latter starts to slightly outperform the supervised classifier at larger masses of $X$. The good agreement between the SC and AD for the $SH$ channel is likely due to the fact that the final state of these BSM models is relatively similar. It is expected that the autoencoder may start to outperform the SC for events that have a substantial difference in the final state, such as the presence of additional identified particles that are not part of the Higgs decays. The anomaly-detection technique has not been used in di-Higgs searches. However, it offers a model-agnostic approach that generalizes more effectively across multiple BSM scenarios compared to a supervised classifier in di-Higgs analyses targeting high-mass resonances.
The Monte Carlo data and the analysis code for this paper are available from [17] and [18].
@@ 5 Acknowledgments
The submitted manuscript has been created by UChicago Argonne, LLC, Operator of Argonne National Laboratory (“Argonne”). Argonne, a U.S. Department of Energy Office of Science laboratory, is operated under Contract No. DE-AC02-06CH11357. The U.S. Government retains for itself, and others acting on its behalf, a paid-up nonexclusive, irrevocable worldwide license in said article to reproduce, prepare derivative works, distribute copies to the public, and perform publicly and display publicly, by or on behalf of the Government. The Department of Energy will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan. http://energy.gov/downloads/doe-public-access-plan. Argonne National Laboratory’s work was funded by the U.S. Department of Energy, Office of High Energy Physics (DOE OHEP) under contract DE-AC02-06CH11357. The Askaryan Calorimeter Experiment was supported by the US DOE OHEP under Award Numbers DE-SC0009937, DE-SC0010504, and DE-AC02-76SF0051. WI is supported by DE-SC0017647. We gratefully acknowledge the computing resources provided by the Laboratory Computing Resource Center at Argonne National Laboratory.

### v2

@@ 4 Conclusion
In summary, we tested two fundamentally different ML approaches: one, a supervised classifier that requires both SM background and BSM signal events as input, and the other, an autoencoder-based anomaly detection approach that only requires SM events for training. The benchmarks were based on models with $X\rightarrow HH$ and $X\rightarrow SH$ boson decays, with the final state involving two leptons and two $b$-jets. The same input variables were used in both ML approaches, and the architecture of the hidden layers was identical. Therefore, these studies provide a compelling comparison of two distinct approaches that use the same inputs. Such a comparison of very different methods, using the same input variables, has been performed for the first time. Note that the anomaly-detection technique has not been used so far in di-Higgs searches. The Monte Carlo data and the analysis code for this paper are available from [Chekanov:2014fga] and [github].
We conclude that both ML techniques were able to improve the significance of signal observations. The event classifier outperforms the autoencoder for the BSM events used in training, especially for low masses of the scalar $X$. This is unsurprising for the $HH$ signals, which were used to train the SC. However, it was unexpected to observe that the classifier works quite well even for the $SH$ signal events, which were not used in training. Its performance is comparable to that of the autoencoder. The latter begins to slightly outperform the supervised classifier at higher values of the $X$ mass. The good agreement between the SC and AD in the $SH$ channel is likely due to the relatively similar final states predicted by these BSM models.
It is expected that the autoencoder starts to outperform the SC for events that have a substantial difference in the final state, such as the presence of additional identified particles that are not part of the Higgs decays. Thus, the method offers a model-agnostic approach that generalizes more effectively across multiple BSM scenarios compared to a supervised classifier in di-Higgs analyses targeting high-mass resonances.
@@ 5 Acknowledgments
The submitted manuscript has been created by UChicago Argonne, LLC, Operator of Argonne National Laboratory (“Argonne”). Argonne, a U.S. Department of Energy Office of Science laboratory, is operated under Contract No. DE-AC02-06CH11357. Argonne National Laboratory’s work was funded by the U.S. Department of Energy, Office of High Energy Physics (DOE OHEP) under contract DE-AC02-06CH11357. We gratefully acknowledge the computing resources provided by the Laboratory Computing Resource Center at Argonne National Laboratory.